x^2=25,other than trial and error,how does one solve for x
Answer by stanbon(75887)

x^2=25
x^2-25=0
(x-5)(x+5)=0
x-5 = 0 or x+5 = 0
x=5 or x=-5
Cheers,
Stan H.